OpenStudy (anonymous):

What is angle A= 82.81924 in degrees and minutes?

OpenStudy (anonymous):

@amistre64 @mathmale

OpenStudy (amistre64):

what does 82.81924 represent, radians or degrees?

OpenStudy (anonymous):

degrees

OpenStudy (amistre64):

then only one part should be giving you some issue, which part would that be?

OpenStudy (anonymous):

I was solving a triangle Angle A, and I used cos A to find that angle... I just forgot how to find the seconds in 82.81924 degrees

OpenStudy (amistre64):

your post says minutes. but its just a conversion ratio ... 60 minutes = 1 degree , multiply both sides by that .81924

OpenStudy (amistre64):

60*60 seconds = 1 degree if you want to find seconds

OpenStudy (anonymous):

Sorry I meant seconds. I'm confused :/ how would I do that 6o times what??

OpenStudy (amistre64):

1 degree = 60 minutes 1 minute = 60 seconds ; *60 gives us 60 minutes = 60*60 seconds therefore 1 degree = 60*60 seconds

OpenStudy (anonymous):

Okay I did mean minutes... I'm all confused lol but I get it now, thanks for the help!

OpenStudy (amistre64):

converting to minutes ... 1 degree = 60 minutes ; * .81924 .81924 degress = 49.1544 minutes

OpenStudy (amistre64):

since 1 minute = 60 seconds, lets *.1544 to get .1544 minutes = 9.264 seconds

OpenStudy (amistre64):

its just a matter of playing with the conversion ratios